OpenStudy (anonymous):

integral lnx^48/x*dx evaluate the indefinite integral?

OpenStudy (lgbasallote):

\[\huge \int \frac{ln (x^{48})}{x}dx\] is that the question?

OpenStudy (lgbasallote):

or is it \[\huge \int \frac{(\ln x)^{48}}{x}dx\]

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\ln ^{48}/x*dx\]

OpenStudy (anonymous):

yes the second one

OpenStudy (lgbasallote):

ahh that's easier let u = ln x du = 1/x dx so the integral becomes \[\huge \implies \int u^{48} du\]

OpenStudy (lgbasallote):

do you get that?

OpenStudy (anonymous):

yes i got that far then did not understand how to integrate

OpenStudy (lgbasallote):

do you remember this rule? \[\huge \int u^n = \frac{u^{n+1}}{n+1}\]

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

it's the power formula...anyway use that

OpenStudy (anonymous):

ok so you get\[u^{49}/49\]

OpenStudy (lgbasallote):

correct. now turn u back into ln x

OpenStudy (anonymous):

ok and thats it right lnx^49/49

OpenStudy (lgbasallote):

yup that's it \[\huge \frac{(\ln x)^{49}}{49}\]

OpenStudy (anonymous):

ok thank you

OpenStudy (lgbasallote):

wellome ^_^